# A short proof of the Hilton-Milner Theorem

### v2

@@ Abstract.
Work of the first author is partially supported by the Israel Science Foundation grant ISF-2480/20. Work of the second author is supported in part by the Slovenian Research Agency (research program P1-0285 and research projects J1-9108, J1-2451, J1-3003, and J1-50000).
@@
Let $k\leq n/2$. If $\mathcal{F}$ is a family of pairwise-intersecting $k$-element subsets of $[n]$, where $\bigcap_{F\in\mathcal{F}}F=\emptyset$, then ||($\left|\mathcal{F}\right|\leq{n-1\choose k-1}-{n-1-k\choose k-1}+1$.
In the current article, we will show Theorem 1 to follow quickly from the following theorem, which we believe to be of some independent interest. Two set systems $\mathcal{A}$ and $\mathcal{B}$ are cross-intersecting if for every $A\in\mathcal{A},B\in\mathcal{B}$, the intersection $A\cap B$ is nonempty. The shadow $\partial B$ of a $k$-element set $B$ consists of all the $(k-1)$-element subsets of $B$; the shadow of a uniform set family is the union of the shadows of its constituent sets.
@@
Let $2k-1\leq n$, let $\mathcal{A}$ be a family of $(k-1)$-element subsets of $[n]$, and let $\mathcal{B}$ be a family of $k$-element subsets of $[n]$. If $\mathcal{A},\mathcal{B}$ are cross-intersecting, $\mathcal{B}$ is nonempty, and $\partial\mathcal{B}\subseteq\mathcal{A}$, then
@@
Note that the bound of Theorem 1 is attained with a single $k$-element set $B$ that does not contain $1$, together with all the $k$-element sets that contain $1$ and intersect $B$; the bound of Theorem 2 is attained with a single $k$-element set and all $(k-1)$-element sets that intersect it.
Our proof of Theorem 1 may be viewed as injective. Other recent proofs of Theorem 1 were given in [4, 10], but instead of relying on a simple cross-intersecting type theorem, both of these proofs rely on a certain “partial complement” operation. In somewhat older work [6] (see also [3]), Frankl and Tokushige gave a proof of Hilton-Milner from a different cross-intersection theorem, but the proof is less elementary than that of Theorem 2, requiring the Schützenberger-Kruskal-Katona Theorem. A recent preprint of Wu, Li, Feng, Liu and Yu [12] gives a proof based on still another cross-intersecting theorem, but the existing proofs of this underlying result also seem to be somewhat more difficult than our approach.
@@
If $n=2k-1$, then the upper bound is (${n\choose k-1}$, and the result follows by noticing that if a $(k-1)$-element set is in $\mathcal{A}$, then its complement cannot be in $\mathcal{B}$ (and vice-versa).
@@
As $\mathcal{A}$ and $\mathcal{B}$ are shifted, so $\mathcal{A}(\neg n)$ and ($\mathcal{B}(\neg n)$ are nonempty, and hence by induction
@@
$\left|\mathcal{A}(n)\right|+\left|\mathcal{B}(n)\right|$ ($\leq{n-1\choose k-2}-{n-1-(k-1)\choose k-2}+1$
($\leq{n-1\choose k-2}-{n-1-k\choose k-2}.$
@@
If $\mathcal{F}$ is a pairwise-intersecting family of $k$-element subsets of $[n]$ with $\bigcap_{F\in\mathcal{F}}F=\emptyset$, then there is a shifted family $\mathcal{F}^{\prime}$ satisfying the same properties and with $\left|\mathcal{F}^{\prime}\right|\geq\left|\mathcal{F}\right|$.
@@
$\mathcal{B}=$$\,\{F$ $:F\in\mathcal{F}\text{ with }1\notin\mathcal{F}\}$
@@
Otherwise, some $\operatorname{Shift}_{i_{0}\leftarrow j_{0}}$ results in a system where every set contains $i_{0}$. Thus, before this step, we have a system $\mathcal{F}$ where every set contains either $i_{0}$ or $j_{0}$. Relabel $i_{0}$ to $1$ and $j_{0}$ to $2$, and continue applying $\operatorname{Shift}_{i\leftarrow j}$ operations over all $3\leq i<j$. Thus, after these additional shift operations, we have $\left\{1,3,\dots,k+1\right\}$ and $\left\{2,3,\dots,k+1\right\}$ in the system. Without loss of generality, we also have all $k$-element subsets containing $\left\{1,2\right\}$; otherwise, add them. Thus, we have $\partial\left\{1,\dots,k+1\right\}$ contained in our system. As $\partial\left\{1,\dots,k+1\right\}$ has empty intersection and is preserved under all further shift operations (over $1\leq i<j$), the result follows. ∎
@@ Discussion
This approach also gives a unified proof of the well-known Erdős-Ko-Rado Theorem. More concretely, if we relax the hypothesis of Theorem 2 to allow $\mathcal{B}$ to be empty, then the corresponding bound is $\left|\mathcal{A}\right|+\left|\mathcal{B}\right|\leq{n-1\choose k-1}$. Erdős-Ko-Rado now follows from replacing Theorem 2 with the relaxed cross-intersection theorem in the proof of Theorem 1. Indeed, the proof is similar and only slightly more complicated than the standard inductive proof of Erdős-Ko-Rado for shifted systems.
@@
A main difficulty in proofs of Hilton-Milner and/or Erdős-Ko-Rado type results is relating systems of $(k-1)$-element subsets to systems of $k$-element subsets. Our approach handles this with the shadow containment condition of Theorem 2.
@@
Our motivation here comes partly from combinatorial algebraic topology. In particular, the simplicial complex generated by a shifted family of $k$-element sets has homology with generators in $\mathcal{B}$ (using notation as in the proof of Theorem 1). Thus, Lemma 3 transforms the combinatorial property of empty intersection into a homological property. Kalai comments on similar connections between intersection theorems and homology in [11, 6.4].

### v4

@@ Abstract.
Work of the first author is partially supported by the Israel Science Foundation grant ISF-2480/20 and the AARMS postdoctoral fellowship. Work of the second author is supported in part by the Slovenian Research Agency (research program P1-0285 and research projects J1-9108, J1-2451, J1-3003, and J1-50000).
@@
Let $k\leq n/2$. If $\mathcal{F}$ is a family of pairwise-intersecting $k$-element subsets of $[n]$, where $\bigcap_{F\in\mathcal{F}}F=\emptyset$, then $\left|\mathcal{F}\right|\leq{n-1\choose k-1}-{n-1-k\choose k-1}+1$.
In the current article, we will show Theorem 1 to follow quickly from the following theorem, which we believe to be of some independent interest. Two set systems $\mathcal{A}$ and $\mathcal{B}$ are cross-intersecting if for every $A\in\mathcal{A},B\in\mathcal{B}$, the intersection $A\cap B$ is nonempty. The shadow $\partial B$ of a $k$-element set $B$ consists of all the (k1)(k-1)-element subsets of $B$; the shadow of a uniform set family is the union of the shadows of its constituent sets, thus consists of all (k1)(k-1)-element subsets of constituent sets.
@@
Let $2k-1\leq n$, let $\mathcal{A}$ be a family of (k1)(k-1)-element subsets of $[n]$, and let $\mathcal{B}$ be a family of $k$-element subsets of $[n]$. If $\mathcal{A},\mathcal{B}$ are cross-intersecting, $\mathcal{B}$ is nonempty, and $\partial\mathcal{B}\subseteq\mathcal{A}$, then
@@
Note that the bound of Theorem 1 is attained with a single $k$-element set $B$ that does not contain $1$, together with all the $k$-element sets that contain $1$ and intersect $B$; the bound of Theorem 2 is attained with a single $k$-element set and all (k1)(k-1)-element sets that intersect it.
Our proof of Theorem 1 may be viewed as injective. Other recent proofs of Theorem 1 were given in [4, 10], but instead of relying on a simple cross-intersecting type theorem, both of these proofs rely on a certain “partial complement” operation. In somewhat older work [6] (see also [3]), Frankl and Tokushige gave a proof of Hilton-Milner from a different cross-intersection theorem, but the proof is less elementary than that of Theorem 2, requiring the Schützenberger-Kruskal-Katona Theorem. A recent preprint of Wu, Li, Feng, Liu and Yu [12] (now published) gives a proof based on still another cross-intersecting theorem, but the existing proofs of this underlying result also seem to be somewhat more difficult than our approach.
@@
If n=2k1n=2k-1, then the upper bound is $n\choose k-1$}, and the result follows by noticing that if a (k1)(k-1)-element set is in $\mathcal{A}$, then its complement cannot be in $\mathcal{B}$ (and vice-versa).
@@
As $\mathcal{A}$ and $\mathcal{B}$ are shifted, so $\mathcal{A}(\neg n)$ and $\mathcal{B}(\neg n)$ are nonempty, and hence by induction
@@
$\left|\mathcal{A}(n)\right|+\left|\mathcal{B}(n)\right|$ $\leq{n-1\choose k-2}-{n-1-(k-1)\choose k-2}+1$
(2) $\leq{n-1\choose k-2}-{n-1-k\choose k-2}$.
@@
If $\mathcal{F}$ is a pairwise-intersecting family of $k$-element subsets of $[n]$ with $\bigcap_{F\in\mathcal{F}}F=\emptyset$, then there is a shifted family $\mathcal{F}^{\prime}$ satisfying the same properties and with $|\left|\mathcal{F}^{\prime}\right|\geq\left$$\mathcal{F}\right|$.
@@
$\mathcal{B}=$ $\,\{F$ $:F\in\mathcal{F}\text{ with }1\notin\mathcal{F}\}$ .
@@
Otherwise, some $\operatorname{Shift}_{i_{0}\leftarrow j_{0}}$ results in a system where every set contains $i_{0}$. Thus, before this step, we have a system $\mathcal{F}$ where every set contains either $i_{0}$ or $j_{0}$. Relabel $i_{0}$ to $1$ and $j_{0}$ to $2$, and continue applying $\operatorname{Shift}_{i\leftarrow j}$ operations over all $3\leq i<j$. Thus, after these additional shift operations, we have $\left\{1,3,\dots,k+1\right\}$ and $\left\{2,3,\dots,k+1\right\}$ in the system. Without loss of generality (since every set in $\mathcal{F}$ contains $1$ or $2$), we also have all $k$-element subsets containing $\left\{1,2\right\}$; otherwise, add them. Thus, we have $\partial\left\{1,\dots,k+1\right\}$ contained in our system. As $\partial\left\{1,\dots,k+1\right\}$ has empty intersection and is preserved under all further shift operations (over $1\leq i<j$), the result follows. ∎
@@ Discussion
A main difficulty in proofs of Hilton-Milner and/or Erdős-Ko-Rado type results is relating systems of (k1)(k-1)-element subsets to systems of $k$-element subsets. Our approach handles this with the shadow containment condition of Theorem 2.
@@
Our motivation here comes partly from combinatorial algebraic topology. In particular, the simplicial complex generated by a shifted family of $k$-element sets has homology with generators in $\mathcal{B}$ (using notation as in the proof of Theorem 1). Thus, Lemma 3 transforms the combinatorial property of empty intersection into a homological property. Kalai comments on similar connections between intersection theorems and homology in [11, Section 6.4].
The approach also gives a unified proof of the well-known Erdős-Ko-Rado Theorem. More concretely, if we relax the hypothesis of Theorem 2 to allow $\mathcal{B}$ to be empty, then the corresponding bound is $\left|\mathcal{A}\right|+\left|\mathcal{B}\right|\leq{n-1\choose k-1}$. Erdős-Ko-Rado now follows from replacing Theorem 2 with the relaxed cross-intersection theorem in the proof of Theorem 1. The proof is similar to (and only slightly more complicated than) the standard inductive proof of Erdős-Ko-Rado for shifted systems.
The approach also recovers uniqueness of the largest family for Theorem 1 when $n/2>k\geq 4$. Here, we strengthen the hypothesis of Theorem 2 to require $\mathcal{B}$ to have at least two elements. We discuss the details in the following section.
Uniqueness of the Hilton-Milner family
As mentioned in the discussion, the same techniques give uniqueness of the maximum family in Theorem 1. We prove:
Theorem 5.
In the situation of Theorem 1, if $4\leq k<n/2$ and $\left|\mathcal{F}\right|$ achieves the upper bound, then there is some $k$-set $B$ and $i\notin B$ so that $\mathcal{F}$ consists of $B$ together with all $k$-sets that both contain $i$ and intersect $B$.
We require $k\geq 4$ in order to avoid some technicalities. In particular, there is another family achieving the bound for $k=3$. See [10] for more details and a different argument.
As in the proof of Theorem 1, we reduce to a shifted family, and prove for a shifted family.
The proof for a shifted family requires a completely straightforward modification of Theorem 2. We obviously require $k\geq 4.$ We also strengthen the hypothesis to require $\left|\mathcal{B}\right|\geq 2$, replacing the condition that $\left|\mathcal{B}\right|\geq 1$; with the strengthened hypothesis, the inequality is strict. Then in the proof, we may have $\left|\mathcal{B}(n)\right|$ empty or nonempty. If empty, then since $\mathcal{B}$ has at least two elements, so $\mathcal{A}(n)$ is strictly smaller than the given bound. If nonempty, then the bound in (2) is already strict so long as $k\geq 4$. In either case, the induction step yields a strict inequality.
Theorem 5 follows for shifted families by applying the variant of Theorem 2 with $\left|\mathcal{B}\right|\geq 2$ to the same families as in the proof of Theorem 1.
It remains only to reduce to shifted families. This reduction requires a bit of care. We did not find the following lemma in the literature, although we believe it to be known to experts in the field.
Lemma 6.
Let $\mathcal{F}$ be a family of pairwise-intersecting $k$-element subsets of $[n]$ with the additional property that for any $F_{0}\in\mathcal{F}$, the intersection $\bigcap_{\mathcal{F}\setminus\{F_{0}\}}F$ is empty. Then there is a shifted family $\mathcal{F}^{\prime}$ satisfying the same properties and with $|\left|\mathcal{F}^{\prime}\right|\geq\left$$\mathcal{F}\right|$.
Proof..
By the standard family, we mean the shifted family with $A=\left\{2,\dots,k+1\right\}$, $A^{\prime}=\left\{2,\dots,k,k+2\right\}$, and all $k$-element sets that both contain $1$ and intersect $A$ and $A^{\prime}$. It is obvious that the standard family is at least as large as any family where all but two sets contain $1$.
Given $\mathcal{F}$, we perform a sequence of shifts. If these terminate in a shifted family with the desired properties, then we are done. Otherwise, an operation results in a family without the additional property. Stopping just before this operation and relabeling elements, we have a family containing sets with $1$ and not $2$, with $2$ and not $1$, with both $1$ and $2$, and possibly the set $B=\{3,\dots,k+2\}$.
We may assume without loss of generality that we have all sets containing both $1,2$ and intersecting with $B$. Since these sets do not have any common intersection other than $1,2$, the operations $\operatorname{Shift}_{i\leftarrow j}$ over all $3\leq i<j$ preserve the additional property.
After shifting over $3\leq i<j$, if we have only one set with $1$ and not $2$, or only one set with $2$ and not $1$, then we replace with the standard family. Otherwise, we have in the family $\{a,3,\dots,k+1\}$ and $\{a,3,\dots,k,k+2\}$ for $a=1,2$, along with all sets containing $\{1,2\}$ and intersecting $B$. In particular, the family contains as subfamilies both $\partial\left\{1,\dots,k+1\right\}$ and $\partial\left\{1,\dots,k,k+2\right\}$. Both subfamilies have empty intersection and are preserved under all shift operations, so we can now shift until the system stabilizes. ∎